TIE RATIO MPI Incorporated has $9 billion in assets, and its tax rate is 40%. Its basic earning power (BEP) ratio is 10%, and its return on assets (ROA) is 3%, what is MPIs times-interest-earned (TIE) ratio? Round your answer to two decimal places
Answer #1

Answer - TIE ratio = 2 times

Reason -

ROA = net income / asset
or, 0.03 = net income / 9,000,000,000
or, net income = 270,000,000

BEP = EBIT/ asset
or, 0.10 = EBIT / 9,000,000,000
or, EBIT = 900,000,000

pre tax net income = 270,000,000 ÷ (1 - 0.40) = 450,000,000

Interest = EBIT - pre tax net income
interest = 900,000,000 - 450,000,000 = 450,000,000

Times interest earned ratio = EBIT / interest
= 900,000,000 / 450,000,000
= 2 times